# The Impact of Spectroscopic Redshift Errors on Cosmological Measurements

### v1

@@ 2.3 Combined Errors for Slitless Redshift Surveys
Space missions such as Euclid [13], Roman [38] and CSST (scheduled for launch in 2027) [39] employ slitless spectroscopy to map the 3D distribution of galaxies over wide areas of sky. Early datasets from Euclid (Quick Data Release 1; [34]) are now publicly available, providing initial evaluations of the performance of the redshift measurement. Unlike traditional slit-based spectroscopy that captures individual targets using fibers, slitless spectroscopy disperses the light from all sources in the field simultaneously, enabling efficient and continuous spectroscopic observations without the need for target pre-selection [29, 17]. This approach allows for high sky coverage and dense sampling of galaxies in the redshift range relevant to LSS studies.
However, the slitless design magnifies the redshift errors. Euclid spectroscopic redshifts are derived from spectra provided by the Near Infrared Spectro-Photometer (NISP), which operates with a resolution of $R\simeq 500$ [29]. In comparison, the ground-based DESI survey uses instruments with resolution $R\sim 2000$ to $5500$ [10]. The lower spectral resolution in slitless survey leads to broader spectral lines and consequently increased redshift uncertainty. Additionally, the lack of spatial isolation between sources causes spectral contamination, typically through cross-contamination and self-contamination [17]. Cross-contamination arises from overlapping spectrograms of neighboring sources (redshift interlopers; [23]), while self-contamination results from the degeneracy between spatial and spectral information along the dispersion direction. These combined effects lead to a higher rate of redshift catastrophics in the sample. To address these challenges, we adopt a hypothetical redshift error that combines both redshift uncertainty and catastrophics to simulate the expected observational systematics in slitless surveys. The model assumes larger uncertainty than those typical of ELGs (i.e., LRG-like) and a higher rate of catastrophic errors (5% with $|\Delta v|>1000\,\rm{km}\cdot s^{-1}$). This hypothetical redshift error will help us to assess their influences on cosmological analyses obtained from space slitless surveys.
@@ 3 Contaminated Mocks
$\sigma_{\Delta v}$$w_{\Delta v}$ ($\rm{km}\cdot s^{-1}$) - 85.7 200 92.7 - 85.7
@@
Table 1: Summary of contaminations and mock catalogs used in our study. Mocks include uncontaminated (clean) one and those affected by redshift errors. Redshift uncertainty is modeled using Gaussian (G) or Lorentzian (L) profiles, characterized by $\sigma_{\Delta v}$ or $w_{\Delta v}$ respectively. We consider uncertainties for LRG-like Gaussian (dv-LRG), QSO-like Gaussian (dv-QSO-G) and QSO-like Lorentzian (dv-QSO-L), using the magnitudes reported from DESI EDR repeat observations [26]. Redshift catastrophic failures are based on analyses of DESI DR1 ELG repeat observations [21], with $f_{c}$ being the contamination rate. We consider $f_{c}=1\%$ with ELG-like catastrophic pattern (dv-ELGcatas) in DESI survey. For slitless-like errors (dv-slitless), we assume combined errors with LRG-like uncertainty and $f_{c}=5\%$ ELG-like catastrophics to mimic the expected conditions in Euclid [17]. The corresponding redshift error distributions for each contaminated mock are shown in Figure˜1.
Figure 1: Comparisons of redshift errors distributions. Left: Typical LRG-like and QSO-like redshift uncertainties, modeled using Gaussian (G) or Lorentzian (L). Right: ELG-like catastrophics in DESI survey and hypothetical slitless-like errors for Euclid, which combines LRG-like redshift uncertainty (left part) and $f_{c}=5\%$ ELG-like catastrophic contamination (right part). These redshift error distributions are used to generate the additional velocity dispersion and construcut our contaminated mock catalogs.
@@
We use the Quijote $N$-body simulation suite to construct our mock catalogs [40]. The Quijote fiducial simulations assume a flat $\Lambda\rm{CDM}$ universe, with cosmology consistent with the Planck 2018 result [41]: $h=0.6711,\,\Omega_{\mathrm{m}}=0.3175,\,\Omega_{\mathrm{b}}=0.049,\,\sigma_{8}=0.834,\,n_{s}=0.9624,\,\sum m_{\nu}=0.0\,{\rm{eV}},\,w=-1$. To study the impact of redshift errors on neutrino mass measurement, we also include the ‘Mnu_p’ catalog with $\sum m_{\nu}=0.1,\mathrm{eV}$. In Quijote simulation, each realization contains $512^{3}$ CDM particles with mass $m_{p}\approx 6.5\times 10^{11}\,M_{\odot}$, evolved within a comoving box of volume $1\,h^{-3}\rm{Gpc}^{3}$. To match the redshift range of interest, we use the snapshots at redshift $z=1.0$. A mass selection with $13.1<\log_{10}(M_{h}/(h^{-1}M_{\odot}))<13.5$ is applied to ensure that the truncated halos represent the dominant structures contributing to the galaxy clustering. We use in total 500 realizations to ensure a minimal statistical noise and allow us to robustly assess the potential biases introduced by redshift errors.
@@
The halo catalogs include information such as the Cartesian positions $X,Y,Z$, the corresponding peculiar velocities $\bm{v}_{\rm{pec}}$, and the halo masses, $M_{h}$. Assuming the line-of-sight is along the $x$-direction, we use the $x$-component of the peculiar velocity, $v_{{\rm{pec}},x}$, to transform halo positions from real space to redshift space by $X_{\rm{rsd}}=X_{\rm{real}}+v_{{\rm{pec}},x}\,(1+z)/H(z)$. The clean mock catalogs are built using the peculiar velocity without modification. To build contaminated mocks with redshift errors, we add an additional velocity shift $\Delta v_{\rm{error}}$ to $v_{{\rm{pec}},x}$. This $\Delta v_{\rm{error}}$ simulates the line-of-sight velocity dispersion introduced by redshift errors.
@@
To generate the $\Delta v_{\rm{error}}$ distribution, we use different profiles to simulate redshift errors for different surveys. For redshift uncertainty, we use the magnitudes reported from repeat observations in DESI Early Data Release (EDR) [26]. Specifically, we apply Gaussian profiles (defined in Eq.˜2.1) with $\sigma_{\Delta v}=85.7\,\rm{km}\cdot s^{-1}$ for LRG-like smearing and $\sigma_{\Delta v}=200\,\rm{km}\cdot s^{-1}$ for QSO-like smearing. To simulate the extended tails observed in QSO redshift errors, we also consider a Lorentzian profile (defined in Eq.˜2.2), with $w_{\Delta v}=95.7,\rm{km}\cdot s^{-1}$, which is truncated with $\Delta v_{\rm{smear}}<2000\,\rm{km}\cdot s^{-1}$ to avoid divergence. A summary of the configurations used for the contaminated mocks is in Table˜1, with the comparison of distributions shown in the left panel of Figure˜1. In all cases, we assume no systematic redshift offset (i.e., $\mu=0$ and $p=0$), consistent with observational findings that show minimal bias in the mean redshift determination. The redshift smearing effects for ELG are relatively small and therefore not included in this redshift uncertainty study [21].
@@
For redshift catastrophics, we adopt a model based on DESI DR1 ELG repeat observations [21]. The velocity shifts caused by catastrophics follow a symmetric distribution centered at zero, with each half modeled by the profile $\mathcal{LN}(\Delta\hat{v},|,0.64,0.25^{2},6.62)$ defined in Eq.˜2.3. We set $f_{c}=1\%$ to represent a ELG-like catastrophics, corresponding to the upper limit of contamination rate in eBOSS and DESI galaxies. As discussed in Section˜2.3, to mimic redshift errors expected in slitless surveys, we consider a more pessimistic scenario combining LRG-like Gaussian smearing ($\sigma_{\Delta v}=85.7\,\rm{km}\cdot s^{-1}$) with $f_{c}=5\%$ ELG-like catastrophics. The contaminated mocks are summarized in Table˜1 and the distributions of ELG-like catastrophics and sliltless-like errors are shown in the right panel of Figure˜1.
@@ 4.1 Power Spectrum Estimator
We estimate the power spectrum of mock catalogs using the weighted galaxy density field with a convenient normalization [42]:
$F(\textbf{r})=\frac{w(\textbf{r})\left[n_{g}(\textbf{r})-\alpha n_{s}(\textbf{r})\right]}{A^{1/2}}$ (4.1)
where $w(\textbf{r})$ is the weight function. $n_{g}$ and $n_{s}$ are the observed number densities of galaxy and random catalogs, respectively. The factor $\alpha=N_{g}/N_{s}$ normalizes the contribution from the random, and the normalization factor $A$ is defined with the expected mean space density of galaxies $\bar{n}(\textbf{r})$ as $A=\int d\textbf{r}\left[\bar{n}(\textbf{r})w(\textbf{r})\right]^{2}$. The estimator of power spectrum multipoles $P_{\ell}$ is given by [42, 43]:
$P_{\ell}(\vec{k})=(2\ell+1)\int\frac{d\Omega_{k}}{4\pi}\left[\int d\textbf{r}_{1}\int d\textbf{r}_{2}F(\textbf{r}_{1})F(\textbf{r}_{2})\times e^{i\vec{k}\cdot(\textbf{r}_{1}-\textbf{r}_{2})}\mathcal{L}_{\ell}(\hat{\textbf{k}}\cdot\hat{\textbf{r}}_{h})-\rm{SN}_{\ell}\right]$ (4.2)
@@ 4.2.1 Effective-field theory of LSS
Effective field theory (EFT) has become a powerful tool to model the LSS of the Universe. By studying small-scale nonlinear physics, the EFT of LSS extend the standard perturbation theory to scales beyond the breakdown of the fluid and single-stream approximations [45, 46, 47]. With loop corrections and improved modeling of biased tracers with redshift-space distortions, EFT enables reliable predictions of the galaxy power spectrum multipoles up to $k\approx 0.25\,h\,\rm{Mpc}^{-1}$ in redshift space [48, 49, 50, 51]. In general, the galaxy redshift-space power spectrum in EFT framework is expressed as [8]:
@@ 4.2.2 Full-Shape Fitting Approaches
We adopt two fitting strategies to perform full-shape cosmological inference: ShapeFit compression [56] and Full-Modeling. These approaches have been extensively tested and validated in analyses of the BOSS and DESI surveys [57, 58, 59, 60, 61].
Compression analysis focuses on fitting a set of compressed physical parameters that describe the shape and the anisotropies of the galaxy power spectrum. These include the $f\sigma_{8}$ and the scaling parameters $\alpha_{\parallel}$ and $\alpha_{\perp}$, which are often reparameterized into the isotropic dilation parameter $\alpha_{\mathrm{iso}}\equiv\alpha_{\parallel}^{1/3}\alpha_{\perp}^{2/3}$ and the Alcock-Paczynski (AP) parameter $\alpha_{\mathrm{ap}}\equiv\alpha_{\parallel}\alpha_{\perp}^{-1}$. ShapeFit further introduces two shape-related parameters, $m$ and $n$, to capture information from the early-time transfer function. In this work, we vary only one shape parameter $m$, while keeping $n$ fixed, and we report ShapeFit results relative to a reference cosmology, specifically using $dm\equiv m-1$ and $df\equiv f/f^{\rm{fid}}$.
Another fitting approach, known as Full-Modeling analysis, avoids the compression steps and directly fits cosmological models to the power spectrum signals. In this way, the parameter dependence of the linear power spectrum along and geometrical information are not kept separate. This allows for a direct comparison between theoretical predictions and observational data, leading direct constraints on cosmological parameters. In our baseline fitting procedure, we sample five cosmological parameters: $h$, $\omega_{\rm cdm}$, $\omega_{b}$, $\ln(10^{10}A_{s})$, $n_{s}$. These parameters are directly used in theoretical models, which facilitates the calculation and a direct comparison of cosmological biases introduced by redshift errors.
In the context of redshift error studies, ShapeFit provides a direct view of how redshift errors distort the shape of observed power spectrum, as reflected by their impact on compressed physical parameters. This makes the interpretation of redshift errors effect straightforward. The Full-Modeling approach, on the other hand, gives cosmological constraints, allowing us to quantify the biases introduced by redshift errors in final cosmological parameter constraints.

### v4

@@ 2.3 Combined errors for Slitless redshift surveys
Space missions such as Euclid [euclid_collaboration_euclid_2025-2], Roman [spergel_wide-field_2015] and CSST (scheduled for launch in 2027; [csst_collaboration_introduction_2025]) employ slitless spectroscopy to map the 3D distribution of galaxies over wide areas of sky. Early dataset from Euclid QR1 is now publicly available [euclid_collaboration_euclid_2025] and provides initial evaluations of the performance of space spectroscopic redshift measurement. Unlike traditional ground slit-based spectroscopy that use fibers to capture individual targets, slitless spectroscopy disperses the light from all sources in the field simultaneously, enabling efficient and continuous observations without the need for target pre-selection. This approach allows for high sky coverage and dense sampling of galaxies in the redshift range relevant to LSS studies.
However, the slitless design magnifies the redshift errors. Euclid spectroscopic redshifts are derived from spectra provided by the Near Infrared Spectro-Photometer (NISP; [euclid_collaboration_euclid_2025-4]), which operates with a resolution of $R\simeq 500$ [euclid_collaboration_euclid_2025-3]. In comparison, the ground-based DESI survey uses instruments with resolution $R\sim 2000$ to $5500$ [desi_collaboration_desi_2016]. The lower spectral resolution and slitless spectroscopy pipeline lead to broader spectral lines and consequently increased redshift uncertainty. Additionally, the lack of spatial isolation between sources causes catastrophic redshift errors [euclid_collaboration_euclid_2025-1, euclid_collaboration_euclid_2025-3], typically through cross-contamination and self-contamination. Cross-contamination arises from overlapping spectrograms of neighboring sources, while self-contamination results from the degeneracy between spatial and spectral information along the dispersion direction. These combined effects lead to a higher rate of redshift catastrophic failures in the sample [euclid_collaboration_euclid_2025-5].
To address these challenges, we adopt a hypothetical redshift error that combines both redshift uncertainty and catastrophics to simulate the expected observational systematics in slitless surveys. This model assumes a larger uncertainty than those typical of ELG (i.e., LRG-like, $\sigma_{\Delta v}=85.7\,\rm{km}\cdot s^{-1}$) and a higher rate of catastrophic failures ($f_{c}=5\%$ with $|\Delta v|>1000\,\rm{km}\cdot s^{-1}$). For reference, [euclid_collaboration_euclid_2025-1] reports that the scatter of redshift difference between Euclid and DESI is $\sigma_{dz}=0.8\times 10^{-3}\approx 240\,\rm{km/s}$, and the overall success rate indicates that a catastrophic rate $f_{c}\gtrsim 10\%$ is expected. Therefore, our hypothetical scenario remains relatively optimistic compared to the current expectations from slitless survey.
@@ 3 Contaminated Mocks
Figure 1: Comparisons of redshift error ($\Delta v_{\rm{error}}$) distributions. Left: Typical LRG-like and QSO-like redshift uncertainties, modeled by Gaussian smearing with $\sigma_{\Delta v}$ or Lorentzian smearing with $\mathrm{w}_{\Delta v}$. Right: ELG-like catastrophics with $f_{c}=1\%$ and hypothetical slitless-like errors with LRG-like redshift uncertainty (left part) and $f_{c}=5\%$ catastrophics (right part). These distributions are used to generate the velocity dispersion caused by redshift errors.
@@
$\sigma_{\Delta v}$$\mathrm{w}_{\Delta v}$ ($\rm{km}\cdot s^{-1}$) - 85.7 300 100 - 85.7
@@
Table 1: Summary of contaminations and mock catalogs used in our study. Mocks include uncontaminated (clean) and those affected by redshift errors. We consider redshift uncertainties for LRG-like Gaussian (dv-LRG), QSO-like Gaussian (dv-QSO-G) and QSO-like Lorentzian (dv-QSO-L) profiles. Redshift catastrophic failures are based on analyses of DESI DR1 ELG repeat observations [yu_elg_2025], with $f_{c}$ being the contamination rate. We consider $f_{c}=1\%$ with ELG-like catastrophic pattern (dv-ELGcatas) in DESI survey. For slitless-like errors (dv-slitless), we assume combined errors with LRG-like uncertainty and $f_{c}=5\%$ ELG-like catastrophics to mimic the expected observing conditions in slitless survey like Euclid. The corresponding redshift error distributions applied in contaminated mocks are shown in Figure˜1.
@@
We use the Quijote $N$-body simulation suite to construct our mock catalogs [villaescusa-navarro_quijote_2020]. The Quijote fiducial simulations assume a flat $\Lambda\rm{CDM}$ universe, with cosmology consistent with the Planck 2018 result [planck_collaboration_planck_2020]: $h=0.6711,\,\Omega_{\mathrm{m}}=0.3175,\,\Omega_{\mathrm{b}}=0.049,\,\sigma_{8}=0.834,\,n_{s}=0.9624,\,\sum m_{\nu}=0.0\,{\rm{eV}},\,w=-1$. To study the impact of redshift errors on neutrino mass measurement, we also include the ‘Mnu_p’ catalog with $\sum m_{\nu}=0.1\,\mathrm{eV}$. In Quijote simulation, each realization contains $512^{3}$ CDM particles with mass $m_{p}\approx 6.5\times 10^{11}\,M_{\odot}$, evolved within a comoving box of volume $1\,h^{-3}\rm{Gpc}^{3}$. To match the redshift range of interest, we use the snapshots at redshift $z=1.0$. A mass selection with $13.1<\log_{10}(M_{h}/(h^{-1}M_{\odot}))<13.5$ is applied to ensure that the truncated halos represent the dominant structures contributing to the galaxy clustering. We use 500 realizations from the Quijote  suite to construct mock catalogs contaminated by different types of redshift errors.
@@
The halo catalogs include information such as the Cartesian positions $X,Y,Z$, the corresponding peculiar velocities $\bm{v}_{\rm{pec}}$, and the halo masses, $M_{h}$. Assuming the line-of-sight is along the $x$-direction, we use the $x$-component of the peculiar velocity, $v_{{\rm{pec}},x}$, to transform halo positions from real space to redshift space at redshift $z$: $X_{\rm{rsd}}=X_{\rm{real}}+v_{{\rm{pec}},x}\,(1+z)/H(z)$. The clean mock catalogs are built using this peculiar velocity without modification. To build contaminated mocks with redshift errors, we add an additional velocity shift $\Delta v_{\rm{error}}$ to $v_{{\rm{pec}},x}$. This $\Delta v_{\rm{error}}$ simulates the line-of-sight velocity dispersion introduced by redshift errors.
@@
To generate the $\Delta v_{\rm{error}}$ distribution, we use different profiles to simulate redshift errors for different surveys. For redshift uncertainty, we use the magnitudes reported from eBOSS and DESI Early Data Release (EDR) [zarrouk_clustering_2018, yu_desi_2023]. Specifically, we apply Gaussian profiles in Eq.˜2.1 with $\sigma_{\Delta v}=85.7\,\rm{km}\cdot s^{-1}$ to simulate LRG-like smearing and $\sigma_{\Delta v}=300\,\rm{km}\cdot s^{-1}$ for QSO-like smearing. We also consider a Lorentzian profile in Eq.˜2.2 to mimic the extended tails observed in QSO redshift uncertainty. The Lorentzian profile is parametrized with $\mathrm{w}_{\Delta v}=100\,\rm{km}\cdot s^{-1}$ and truncated by $\Delta v_{\rm{smear}}<3000\,\rm{km}\cdot s^{-1}$. Using these values, we have almost the same RMS of $\Delta v_{\rm{smear}}$ for the QSO Gaussian and Lorentzian smearing. The redshift smearing effects for ELG are small ($\sigma_{\Delta v}\sim 10\,\rm{km}\cdot s^{-1}$; [yu_elg_2025]) and we neglect them in this study. A summary of the configurations for the contaminated mocks is in Table˜1, with the comparison of distributions shown in the left panel of Figure˜1. In all cases, we assume the redshift errors are symmetric to zero (i.e., $\mu=0$ and $p=0$).
@@
For redshift catastrophics, we adopt a model based on DESI DR1 ELG repeat observations [yu_elg_2025]. The velocity shifts caused by catastrophics follow a symmetric distribution centered at zero, with each half modeled by the profile $\mathcal{LN}(\Delta\hat{v}_{c}\,|\,0.64,0.25^{2},6.62)$ defined in Eq.˜2.3. We set $f_{c}=1\%$ to represent ELG-like catastrophics, corresponding to the contamination rate in eBOSS and DESI galaxies. As discussed in section˜2.3, to mimic redshift errors expected in slitless surveys, we consider a scenario combining LRG-like Gaussian smearing ($\sigma_{\Delta v}=85.7\,\rm{km}\cdot s^{-1}$) with $f_{c}=5\%$ ELG-like catastrophics. The contaminated mocks are summarized in Table˜1 and the distributions of ELG-like catastrophics and sliltless-like errors are shown in the right panel of Figure˜1.
@@ 4.1 Power spectrum estimator
We estimate the power spectrum of mock catalogs using the weighted galaxy density field with a convenient normalization [feldman_power_1994]:
$F(\textbf{r})=\frac{w(\textbf{r})\left[n_{g}(\textbf{r})-\alpha n_{s}(\textbf{r})\right]}{A_{g}^{1/2}}$ (4.1)
where $w(\textbf{r})$ is the weight function. $n_{g}$ and $n_{s}$ are the observed number densities of galaxy and random catalogs, respectively. The factor $\alpha=N_{g}/N_{s}$ normalizes the contribution from the random, and the normalization factor $A$ is defined with the expected mean space density of galaxies $\bar{n}(\textbf{r})$ as $A_{g}=\int d\textbf{r}\left[\bar{n}(\textbf{r})w(\textbf{r})\right]^{2}$. The estimator of power spectrum multipoles $P_{\ell}$ is given by [feldman_power_1994, hand_optimal_2017]:
$P_{\ell}(k)=(2\ell+1)\int\frac{d\Omega_{k}}{4\pi}\left[\int d\textbf{r}_{1}\int d\textbf{r}_{2}F(\textbf{r}_{1})F(\textbf{r}_{2})\times e^{i\ \textbf{k}\cdot(\textbf{r}_{1}-\textbf{r}_{2})}\mathcal{L}_{\ell}(\hat{\textbf{k}}\cdot\hat{\textbf{r}}_{h})-\rm{SN}_{\ell}\right]$ (4.2)
@@ 4.2.1 Effective-field theory of LSS
Effective field theory (EFT) has become a powerful tool to model the LSS of the Universe. By studying small-scale nonlinear physics, the EFT of LSS extends the standard perturbation theory to scales beyond the breakdown of the fluid and single-stream approximations [baumann_cosmological_2012, carrasco_effective_2012, vlah_lagrangian_2015]. With loop corrections and improved modeling of biased tracers with redshift-space distortions, EFT enables reliable predictions of the galaxy power spectrum multipoles up to $k\approx 0.25\,h\,\rm{Mpc}^{-1}$ in redshift space [mcdonald_clustering_2009, senatore_redshift_2014, senatore_bias_2015, chen_redshift-space_2021, aviles_redshift_2021]. In general, the galaxy redshift-space power spectrum in EFT framework is expressed as [desi_collaboration_desi_2025-1]:
@@ 4.2.2 Full-Shape fitting approaches
We adopt two fitting strategies to perform full-shape cosmological inference: ShapeFit compression [brieden_shapefit_2021] and Full-Modeling. These approaches have been extensively tested and validated in analyses of the BOSS and DESI surveys [maus_comparison_2023, noriega_comparing_2024, maus_analysis_2024, lai_comparison_2024, desi_collaboration_desi_2025-1].
Compression analysis focuses on fitting a set of compressed physical parameters that describe the shape and the anisotropies of the galaxy power spectrum. These include the scaling parameters $\alpha_{\parallel}$, $\alpha_{\perp}$ and the growth rate $f\sigma_{8}$. The scaling parameters are often reparameterized into the isotropic dilation parameter $\alpha_{\mathrm{iso}}\equiv\alpha_{\parallel}^{1/3}\alpha_{\perp}^{2/3}$ and the Alcock-Paczynski (AP) parameter $\alpha_{\mathrm{ap}}\equiv\alpha_{\parallel}\alpha_{\perp}^{-1}$. ShapeFit further introduces two shape-related parameters, $m$ and $n$, to capture information from the early-time transfer function. In this work, we vary only one shape parameter $m$, while keeping $n$ fixed [brieden_shapefit_2021, desi_collaboration_desi_2025-1]. We report ShapeFit results relative to a reference cosmology, specifically using $dm\equiv m-1$ and $df\equiv f/f^{\rm{fid}}$.
Another fitting approach, known as Full-Modeling analysis, avoids the compression steps and directly fits cosmological models to the power spectrum signals. In this way, the parameter dependence of the linear power spectrum along and geometrical information are not kept separate. This allows for a direct comparison between theoretical predictions and observational data and thus gives direct constraints on parameters. In our baseline fitting procedure, we consider five cosmological parameters: $h$, $\omega_{\rm cdm}$, $\omega_{b}$, $\ln(10^{10}A_{s})$, $n_{s}$. These parameters are directly used in theoretical models, which facilitates the calculation and a direct comparison of cosmological biases introduced by redshift errors.
In the context of redshift error studies, ShapeFit provides a direct view of how redshift errors distort the shape of observed power spectrum, as reflected by their impact on compressed physical parameters. This makes the interpretation of redshift errors effect straightforward. The Full-Modeling approach, on the other hand, gives cosmological constraints, allowing us to quantify the biases introduced by redshift errors in final cosmological constraints.